OpenStudy (anonymous):

Line NO is tangent to the circle at N and the measure of arc LM is 64°. What is the measure of angle MNO? thats my question . just need to see how to figuree it out

OpenStudy (amistre64):

it would help if we had a picture to go by

OpenStudy (amistre64):

is LN a diameter of do we assume its just a general chord?

OpenStudy (amistre64):

i think it would have to be a diameter

OpenStudy (amistre64):

theres a thrm someplace that says the angle that sits on the rim of a circle is half the arc length across from it

OpenStudy (anonymous):

It is, The measure of MNO is 58°.

OpenStudy (amistre64):

since arcLM is 64, that makes that angle = 64/2 = 32

OpenStudy (amistre64):

we know the tangent line is 180; and we have 90 and 32 to eat up

OpenStudy (amistre64):

180 - 90 - 32 = whats left

OpenStudy (anonymous):

:-) Thank you 58

OpenStudy (amistre64):

yep